OpenStudy (anonymous):

Some more calculus help... I need it worked out and explanations, so I can do it on my own please!!!

OpenStudy (anonymous):

A car is traveling at 50 mi/h when the brakes are fully applied, producing a constant deceleration of 44 ft/s2. What is the distance covered before the car comes to a stop? (Give your answers correct to one decimal place.)

OpenStudy (anonymous):

look at example 7 and nice worked out problem almost identical to the one you have. you can ignore the first part about converting units because your problem is already in feet per second http://zimmer.csufresno.edu/~kbyler/archive/2008-2/m75B/hw/75B-HW9.pdf

OpenStudy (anonymous):

50*5280/3600 = 220/3 ft/s therefore v(t)=44t+220/3